20X-92=87
I realize that the answer is -0.25 but can some please break it down, if figured it but not sure how I did it. Help! Thanks

Answer by checkley71(8403)

20X-92=87
20X=87+92
20X=179
X=179/20
X=8.95
WHY DO YOU SAY THE ANSWER IS -.25?
ONLY IF THE PROBLEM IS 20X-92+87
THEN
20X-92+87
20X+5
5(4X+1)
OR
20X-92+87=0
20X-92+87=0
20X=-5
X=-5/20
YOU NEED TO CHECK YOUR ENTRY TO ENSURE IT IS CORRECT OR WE CAN'T ALWAYS HELP YOU SOLVE THEM!!!!!!!!! :)
X=-.25